OpenStudy (anonymous):

How would you integrate (x^2)( sqrt of (1+(x^3)) )?

OpenStudy (anonymous):

oh, i meant that x^2 is multiplied to the the second term.

OpenStudy (anonymous):

um du= 3x^2

zepdrix (zepdrix):

Oh not division? :O my bad

OpenStudy (anonymous):

it's okay; i appreciate your help in typing out the steps

OpenStudy (anonymous):

:)

zepdrix (zepdrix):

\[\Large\bf\sf \int\limits \sqrt{\color{royalblue}{1+x^3}}\left(\color{orangered}{x^2\;dx}\right)\]

zepdrix (zepdrix):

\[\Large\bf\sf \color{royalblue}{u=1+x^3}\]\[\Large\bf\sf du=3\color{orangered}{x^2\;dx}\]

zepdrix (zepdrix):

Ok good, see how our du ALMOST matches up with the stuff in the brackets?

OpenStudy (anonymous):

yes.

OpenStudy (shamil98):

\[\large \int\limits_{}^{} x^2 \sqrt{1+x^3}~dx\] \[\large \int\limits_{}^{} \sqrt{1+x^3}~x^2dx\] \[u = 1 + x^3\] \[du = 3x^2 dx\] \[\frac{ 1 }{ 3 } du = x^2 dx\] Set it up with u-sub now.

zepdrix (zepdrix):

...

OpenStudy (anonymous):

is it \[\sqrt{1+x ^{3}} (1/3)du\]

zepdrix (zepdrix):

Before integrating?

zepdrix (zepdrix):

You replaced the differential dx correctly. We need to replace the x also though,\[\Large\bf\sf \int\limits\limits \sqrt{\color{royalblue}{1+x^3}}\left(\color{orangered}{x^2\;dx}\right)\quad =\quad \int\limits\limits \sqrt{\color{royalblue}{u}}\left(\color{orangered}{\frac{1}{3}\;du}\right)\]

OpenStudy (shamil98):

you can factor out the constant (1/3)

zepdrix (zepdrix):

\[\Large\bf\sf \frac{1}{3}\int\limits u^{1/2}\;du\]Just apply the normal Power Rule from there! :)

OpenStudy (anonymous):

zepdrix, where did you get the blue sqrt of u from? sorry if it's a dumb question

zepdrix (zepdrix):

That was our initial thing that we chose to be u,\[\Large\bf\sf \color{royalblue}{u=1+x^3}\]The substitution we chose^

OpenStudy (anonymous):

ohh

OpenStudy (anonymous):

okay i will apply the upper limit of 2 and lower limit of 0 now

OpenStudy (anonymous):

after i integrate.

OpenStudy (anonymous):

thank you

zepdrix (zepdrix):

np. Keep in mind that those limits of integration apply to x! Not u. After integrating, you will need to `undo` your substitution before you plug those values in.

OpenStudy (anonymous):

oh how do i undo it?

OpenStudy (anonymous):

do i replace the u with 1+x^3?

zepdrix (zepdrix):

Yes, after you integrate, before plugging in your 0 and 2. Put 1+x^3 back in place of your u.

OpenStudy (anonymous):

thx

OpenStudy (anonymous):

um how would i use u substitution for How do you integrate [(cosx)^4] (sinx)?

OpenStudy (anonymous):

i tried u=cosx but got the wrong answer in the end

OpenStudy (anonymous):

nvm i got it